OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!! Edward is playing a game where he draws cards with integers on them from a deck of cards. If the integer is positive he moves forward that many steps, if the integer is negative he moves back that many steps. Edward drew the following cards: 3, -3, -4, 5, 10, -7. Where does Edward end up in relation to where he started? A) 4 steps behind where he started. B) 4 steps in front of where he started. C) 32 steps behind where he started. D) 32 steps in front of where he started.

OpenStudy (mvpriest6):

what do you think it is

OpenStudy (mvpriest6):

@CandyMuffin

OpenStudy (superflytyguy):

4 steps in front of where he started.

OpenStudy (superflytyguy):

3 -3 = 0 0 - 4= -4 -4 +5 = 1 1 + 10 = 11 11 - 7 = 4

OpenStudy (mvpriest6):

just use a calculator and add all of them together then u have your answer which equals four and then since its positive we are ahead of where we started

OpenStudy (anonymous):

yes 4 steps in front